OpenStudy (anonymous):

Can someone double check my answer please? Thank You! I got 6 meters. How far will a ball roll across the grass if a friction force of 20N does 120J of work bringing it to a stop?

OpenStudy (anonymous):

@dan815 @AlexandervonHumboldt2 @tkhunny @Ashleyisakitty @Preetha @imqwerty @inkyvoyd @Elsa213 @Daniellelovee @Gokuporter @NotTim @AloneS

OpenStudy (quantummechanics):

You have applied \[W = Fd \cos \theta\] the angle theta is 0 degrees so we can say cos theta = 1 which simplifies to \[W = Fd \implies 120J = (20N)d \implies d = \frac{ 120J }{ 20N } = 6 \frac{ J }{ N } = 6m\]

OpenStudy (priyar):

yes your answer is correct!